# Sample-Efficient Reinforcement Learning of Koopman eNMPC

### v1

@@ 1 Introduction
Alternatively, data-driven (eN)MPCs can be trained using reinforcement learning (RL) methods (see Fig. 1(a)). Compared to SI, such “end-to-end” learning approaches can lead to improved performance in specific control tasks (see, e.g., Chen et al., (2019); Gros and Zanon, (2019); Mayfrank et al., 2024b ; Mayfrank et al., 2024a ).
@@
To this end, a differentiable (e)NMPC policy is constructed, wherein the learnable parameters can be parameters of the dynamic model (see Fig. 1(b), e.g., Mayfrank et al., 2024b ; Mayfrank et al., 2024a ), or other parameters that appear in the objective function or constraints of the (e)NMPC (Gros and Zanon, (2019); Brandner et al., (2023); Brandner and Lucia, (2024)), e.g., the state bounds (see Fig. 1(c)). The latter approach optimizes the policy by compensating for model errors, e.g., via bounds adaptation, and thus needs to learn only few (interpretable) parameters which may lead to better convergence, compared to RL-based training/refinement of the dynamic model itself. However, even positing good training convergence, the model-free [^1] Due to the multitude of ways in which models can be used in RL, the distinction between model-free RL, model-based RL, and other control approaches is not consistent across the literature. This distinction might be especially confusing in the context of the present work since we study the training of eNMPC policies, i.e., inherently model-based policies. However, the distinction between model-free and model-based, which is most relevant to our work, is about whether the training algorithm uses an additional learned model of the environment to train the policy. Please refer to Section 2.1 for the exact definitions that we use in this work. RL algorithms that have so far been used for RL-based training of (eN)MPCs remain notoriously sample inefficient, essentially rendering them inapplicable to domains where interacting with the physical environment (sampling) is expensive (Gopaluni et al., (2020)), e.g., in industrial chemical process control applications.
@@ 2.3 Physics-informed MBPO of Koopman models for control
Figure 3: Using MBPO to train a task-optimal Koopman (e)NMPC controller. (a) The training algorithm. The three steps are executed in a loop until a stopping criterion is reached. (b) The automatically differentiable Koopman (e)NMPC.
First, the Koopman (e)NMPC policy interacts with the environment for a predefined number of steps to gather data $\mathcal{D}$ about the system dynamics. To ensure exploration, we sample the (otherwise deterministic) action $\bm{u}_{t}$ from a normal distribution $\bm{u}_{t}\sim\bm{\mathcal{N}}(\bm{u}^{*}_{t},\bm{\sigma}^{2})$. We assume that the reward function of the environment is known. Thus, rewards $r_{t}$ do not need to be recorded in $\mathcal{D}$.
Second, an ensemble of $n$ data-driven dynamic models, i.e., neural networks (NNs), is learned based on $\mathcal{D}$ via SI to approximate the dynamics of the environment (c.f. Eq. (1)). If (incomplete) physics knowledge is available, it is possible to train PINNs (Raissi et al., 2019b ). Throughout this work, each ensemble member $\bm{NN}_{i,\bm{\omega}_{i}}\forall i\in\{1,2,\dots,n\}$ is a PINN parameterized by $\bm{\omega}_{i}$. Furthermore, we fit the parameters $\bm{\theta}_{\text{K}}$ of the Koopman model to $\mathcal{D}$ via SI. Note that we do not use a physics-informed training method for the Koopman model. In fitting the PINN ensemble and the Koopman model, we follow standard SI practices, such as splitting $\mathcal{D}$ into a training data set and a validation data set used for early stopping and normalizing the inputs and outputs of the models. We refer the reader to Mayfrank et al., 2024b for a detailed description of how Koopman models in the form proposed by Korda and Mezić, (2018) (Eq. (3)) can be trained using system identification.
@@
We emphasize that the behavior of the resulting (e)NMPC is defined by $\bm{\theta}_{\text{K}}$ and $\bm{\theta}_{\text{B}}$ and that both parameter types are optimized in each iteration of the overall algorithm (Fig. 3): In the SI step, $\bm{\theta}_{\text{K}}$ is trained to maximize the agreement of the Koopman model to $\mathcal{D}$, i.e., all data available at the time. Then, the policy optimization step optimizes $\bm{\theta}_{\text{B}}$, i.e., the state bounds, using PPO and simulated interactions with the PINN ensemble.
In the following, we provide three rationales for why training an (e)NMPC policy via RL on an environment learned using SI offers benefits compared to simply learning a Koopman model and using it inside an (e)NMPC: (i) The PINN ensemble captures the epistemic uncertainty of the learned dynamics (Janner et al., (2019)), i.e., the uncertainty that arises from a lack of sufficient amounts of data. Through RL, the (e)NMPC is forced to behave in a way that is robust with respect to every ensemble member. Without RL, a single learned Koopman model would define the (e)NMPC behavior, which increases the risk of policy failures due to epistemic uncertainty. (ii) The representational capacity of the Koopman model that is utilized in the (e)NMPC is limited since the resulting OCPs (Eq. (5)) have to be solved in real-time. Thus, given sufficiently complex system dynamics, the Koopman model might not be able to accurately capture the system dynamics everywhere. In contrast, the representational capacity of the PINN ensemble members has a much higher limit since the ensemble models are not used in online optimization. RL tuning of $\bm{\theta}_{\text{B}}$ can, therefore, help to compensate for the limitations of (e)NMPC that stem from a limited representational capacity of the utilized Koopman model. (iii) Using RL, the (e)NMPC can be tuned to the requirements of a specific control problem. Specifically, by weighting different parts of the reward function of the simulated RL environment, RL offers a way to tune the behavior of the policy, e.g., to prioritize cost savings or constraint satisfaction.
@@ 3.2.1 Model architecture
In the eNMPC, we choose a prediction horizon $t_{f}$ of nine hours, i.e., $\mathrm{T}_{+1}=\{t,\dots,t+9\}$ and $\mathrm{T}=\{t,\dots,t+8\}$ (see Eq. (5)). Analogous to our earlier work on model-free RL of Koopman eNMPCs (Mayfrank et al., 2024b ), given a prediction for the evolution of the electricity prices $\bm{p}_{\text{eNMPC}}=(p_{t})_{t\in\mathrm{T}_{+1}}$, the policy aims to minimize the production cost while satisfying the bounds of the states and the product storage. To that end it first calculates the initial latent state $\bm{z}_{0}$ by passing the initial state $\bm{x}_{0}=(c_{0},T_{0})^{\top}$ through the encoder, i.e., $\bm{z}_{0}=\bm{\psi}_{\bm{\theta}_{\text{K}}}(\bm{x}_{0})$. Then, the following OCP is solved:
@@
For the purpose of data-driven modeling, we rescale the system states and control inputs (see Table 2) linearly so that the lower and upper bound of each variable correspond to $-1.0$ and $1.0$, respectively.
@@ 3.3 Results
Looking at Fig. 6(c), we can see that initially (up to around 1000 steps) $\text{SI}_{\text{Koop}}\text{PIRL}_{\text{Bounds}}$ and $\text{SI}_{\text{Koop}}\text{RL}_{\text{Bounds}}$ are best at avoiding constraint violations. In particular, $\text{SI}_{\text{Koop}}\text{PIRL}_{\text{Bounds}}$ quickly learns to avoid constraint violations. Both variants continue to make relatively steady progress in this regard up until the end of the training runs. The MLP controllers intially cause many constraint violations, however, after around 1000 environment steps they have learned to avoid constraint violations almost perfectly. Towards the end of the training, some of the $\text{RL}_{\text{MLP}}$ controllers seem to loose this capability partially. $\text{SI}_{\text{Koop}}$ struggles with constraint satisfaction across the full length of all training runs.
@@
Figure 8: Closed-loop prediction results. The bounds of each variable (see Tab. 2) are used for scaling to the [-1, 1] range. The red line depicts the true trajectory of $c$ and $T$. Each blue line corresponds to the closed-loop prediction of one ensemble member. Predictions are chained over the full horizon (168 steps) of the episode.
@@ 4 Conclusion
We validate our method using a demand response case study (Mayfrank et al., 2024b ) based on a benchmark CSTR model (Flores-Tlacuahuac and Grossmann, (2006)). The case study involves nonlinear dynamics and hard constraints on system variables; however, due to its small scale it is far less complex than many real-world systems. We compare the performance of our method to that of Koopman eNMPCs trained solely via iterative SI and to NN policies trained via (physics-informed) model-based RL. We find that our method (see 2.3) outperforms all other tested approaches when applied to our case study.

### v2

@@ 1 Introduction
Alternatively, using reinforcement learning (RL) methods, data-driven (eN)MPCs can be trained for optimal performance in specific control tasks (see Fig. 1(a)), which may produce superior control performance compared to SI (see, e.g., Chen et al., (2019); Gros and Zanon, (2019); Mayfrank et al., 2024b ; Mayfrank et al., 2024a ).
@@
To this end, a differentiable (e)NMPC policy is constructed, wherein the learnable parameters can be parameters of the dynamic model (see Fig. 1(b), e.g., Mayfrank et al., 2024b ; Mayfrank et al., 2024a ), or other parameters that appear in the objective function or constraints of the (e)NMPC (Gros and Zanon, (2019); Brandner et al., (2023); Brandner and Lucia, (2024)), e.g., the state bounds (see Fig. 1(c)). The latter approach optimizes the policy by compensating for model errors, e.g., via bounds adaptation. Thereby, RL-based training/refinement of a highly parameterized dynamic model such as an artificial neural network can be avoided by optimizing few (interpretable) parameters, e.g., parameters that modify the state bounds. Therefore, this approach may lead to better convergence, compared to RL-based (re)training of the dynamic model itself. However, even positing good training convergence, the model-free [^1] Due to the multitude of ways in which models can be used in RL, the distinction between model-free RL, model-based RL, and other control approaches is not consistent across the literature. This distinction might be especially confusing in the context of the present work since we study the training of eNMPC policies, i.e., inherently model-based policies. However, the distinction between model-free and model-based, which is most relevant to our work, is about whether the training algorithm uses an additional learned model of the environment to train the policy. Please refer to Section 2.1 for the exact definitions that we use in this work. RL algorithms that have so far been used for RL-based training of (eN)MPCs remain notoriously sample inefficient, essentially rendering them inapplicable to domains where interacting with the physical environment (sampling) is expensive (Gopaluni et al., (2020)), e.g., in industrial chemical process control applications.
@@
In this work, we choose the MBPO algorithm (Janner et al., (2019)), since it is a state-of-the-art Dyna-style RL algorithm. Many model-based RL algorithms exist and it is impossible to know a priori which algorithm will perform best in a specific task (Wang et al., (2019)). While MBPO is one of the most promising algorithms available, our approach should be compatible with any Dyna-style algorithm that does not require a specific policy architecture.
@@ 2.3 Physics-informed MBPO of Koopman models for control
This section describes our general framework for sample-efficient learning of task-optimal Koopman (e)NMPC policies. A more in-depth description of the implementation details of our method when applied to our specific case study is provided in Sec. 3.2.
@@
Figure 3: Using MBPO to train a task-optimal Koopman (e)NMPC controller. (a) The training algorithm. The following three steps are executed in a loop until a stopping criterion is reached: First, the Koopman (e)NMPC interacts with the environment to gather data about the dynamics. Second, all data collected up to the current step is used to fit the Koopman model (parameters $\bm{\theta}_{\text{K}}$) and the PINN ensemble (parameters $\bm{\omega}_{i}\forall i\in\{1,2,\dots,n\}$). Third, a surrogate RL environment is constructed using the NN ensemble and the Koopman (e)NMPC is optimized by tuning the parameters $\bm{\theta}_{\text{B}}$, i.e., the state bounds. (b) The automatically differentiable Koopman (e)NMPC whose behavior is defined by the parameters of the Koopman model ($\bm{\theta}_{\text{K}}$) and the parameters modifying the state bounds ($\bm{\theta}_{\text{B}}$). The parameters are color-coded to match the colors of the corresponding optimization steps in Fig. 3(a).
First, the Koopman (e)NMPC policy interacts with the environment for a predefined number of steps to gather data $\mathcal{D}$ about the system dynamics. To ensure exploration, we sample the (otherwise deterministic) action $\bm{u}_{t}$ from a normal distribution $\bm{u}_{t}\sim\bm{\mathcal{N}}(\bm{u}^{*}_{t},\bm{\sigma}^{2})$. We assume that the reward function of the environment is known. Thus, rewards $r_{t}$ do not need to be recorded in $\mathcal{D}$. In the very first iteration of the overall algorithm the Koopman model is still randomly initialized and the eNMPC outputs will, therefore, not be meaningful. Therefore, in the data sampling step of the first MBPO iteration, we randomly sample the actions $\bm{u}_{t}$ from a uniform distribution over the action space. Likewise, any other controller type, e.g., a PID controller, may be used in the first MBPO iteration, although the quality of the resulting training data will be influenced by how diverse the control actions produced by the controller are.
Second, an ensemble of $n$ data-driven dynamic models, i.e., neural networks (NNs), is learned based on $\mathcal{D}$ via SI to approximate the dynamics of the environment (c.f. Eq. (1)). If (incomplete) physics knowledge is available, it is possible to train PINNs (Raissi et al., 2019b ). Throughout this work, each ensemble member $\bm{NN}_{i,\bm{\omega}_{i}}\forall i\in\{1,2,\dots,n\}$ is a PINN parameterized by $\bm{\omega}_{i}$. Furthermore, we fit the parameters $\bm{\theta}_{\text{K}}$ of the Koopman model to $\mathcal{D}$ via SI. Note that we do not use a physics-informed training method for the Koopman model. Physics-informed training of the Koopman model would, in principle, also be possible. However, the representational capacity of the Koopman model is limited because it is used as part of the real-time eNMPC policy and physics-informed training would necessitate using some of that representational capacity to predict outputs that are not necessary for the eNMPC application, i.e., the a priori unknown physics terms (see Fig. 5). Since the behavior of the overall Koopman eNMPC policy is optimized with respect to the physics-informed NN ensemble anyway (see Fig. 3(a), step three), and to keep our method as simple as possible, we decided against a physics-informed system identification approach for the Koopman model. In fitting the PINN ensemble and the Koopman model, we follow standard SI practices, such as splitting $\mathcal{D}$ into a training data set and a validation data set used for early stopping and normalizing the inputs and outputs of the models. We refer the reader to Mayfrank et al., 2024b for a detailed description of how Koopman models in the form proposed by Korda and Mezić, (2018) (Eq. (3)) can be trained using system identification.
@@
We emphasize that the behavior of the resulting (e)NMPC is defined by $\bm{\theta}_{\text{K}}$ and $\bm{\theta}_{\text{B}}$ and that both parameter types are optimized in each iteration of the overall algorithm (Fig. 3): In the SI step, $\bm{\theta}_{\text{K}}$ is trained to maximize the agreement of the Koopman model to $\mathcal{D}$, i.e., all data available at the time. Then, the policy optimization step optimizes $\bm{\theta}_{\text{B}}$, i.e., the state bounds, using PPO and simulated interactions with the PINN ensemble. Please note that the policy optimization step does not necessarily result in a tightening of the state bounds: Depending on (i) the disagreement between the Koopman model and the PINN ensemble, and (ii) the reward function specified by the user, it may lead to tightened or relaxed bounds. Therefore, if a sensible reward function is specified, the policy optimization step should not be detrimental to the economic performance of the overall policy.
Intuitive justification for our approach:
The policy optimization step optimizes the Koopman (e)NMPC policy using a surrogate RL environment. This environment is based on an ensemble of PINNs learned via SI using the same data as the Koopman model. In the following, we provide three rationales for why such an approach offers benefits compared to simply learning a Koopman model via SI and using it inside an (e)NMPC policy without further RL-based optimization of the policy: (i) The PINN ensemble captures the epistemic uncertainty of the learned dynamics (Janner et al., (2019)), i.e., the uncertainty that arises from insufficient amounts of data. Through RL, the (e)NMPC is forced to behave in a way that is robust with respect to every ensemble member. Without RL, a single learned Koopman model would define the (e)NMPC behavior, which increases the risk of policy failures due to epistemic uncertainty. (ii) The representational capacity of the Koopman model that is utilized in the (e)NMPC is limited since the resulting OCPs (Eq. (5)) have to be solved in real-time. Thus, given sufficiently complex system dynamics, the Koopman model might not be able to accurately capture the system dynamics everywhere. In contrast, the representational capacity of the PINN ensemble members has a much higher limit since the ensemble models are not used in online optimization. RL tuning of $\bm{\theta}_{\text{B}}$ can, therefore, help to compensate for the limitations of (e)NMPC that stem from a limited representational capacity of the utilized Koopman model. (iii) Using RL, the (e)NMPC can be tuned to the requirements of a specific control problem. Specifically, by weighting different parts of the reward function of the simulated RL environment, RL offers a way to tune the behavior of the policy, e.g., to prioritize cost savings or constraint satisfaction.
@@ 3.2.1 Model architecture
RL-based training of (e)NMPC controllers requires solving and differentiating through many OCP instances. Since the number of variables in an OCP grows linearly with the number of time steps, RL training of (e)NMPCs is computationally challenging given long prediction horizons. To balance control performance and computational tractability, we determine an effective eNMPC prediction horizon by repeatedly solving eNMPC problems using the mechanistic CSTR model while varying the prediction horizon. We select a horizon $t_{f}$ of nine hours, as longer horizons do not produce substantial performance gains in the mechanistic eNMPC. Thus, $\mathrm{T}_{+1}=\{t,\dots,t+9\}$ and $\mathrm{T}=\{t,\dots,t+8\}$ (see Eq. (5)). Analogous to our earlier work on model-free RL of Koopman eNMPCs (Mayfrank et al., 2024b ), given a prediction for the evolution of the electricity prices $\bm{p}_{\text{eNMPC}}=(p_{t})_{t\in\mathrm{T}_{+1}}$, the policy aims to minimize the production cost while satisfying the bounds of the states and the product storage. To that end it first calculates the initial latent state $\bm{z}_{0}$ by passing the initial state $\bm{x}_{0}=(c_{0},T_{0})^{\top}$ through the encoder, i.e., $\bm{z}_{0}=\bm{\psi}_{\bm{\theta}_{\text{K}}}(\bm{x}_{0})$. Then, the following OCP is solved:
@@
For the purpose of data-driven modeling, we rescale the system states and control inputs (see Table 2) linearly so that the lower and upper bounds of each variable correspond to $-1.0$ and $1.0$, respectively.
@@ 3.3 Results
Looking at Fig. 6(c), we can see that initially (up to around 1000 steps) $\text{SI}_{\text{Koop}}\text{PIRL}_{\text{Bounds}}$ and $\text{SI}_{\text{Koop}}\text{RL}_{\text{Bounds}}$ are best at avoiding constraint violations. In particular, $\text{SI}_{\text{Koop}}\text{PIRL}_{\text{Bounds}}$ quickly learns to avoid constraint violations. Both variants continue to make relatively steady progress in this regard up until the end of the training runs. The MLP controllers initially cause many constraint violations, however, after around 1000 environment steps they have learned to avoid constraint violations almost perfectly. Towards the end of the training, some of the $\text{RL}_{\text{MLP}}$ controllers seem to loose this capability partially. $\text{SI}_{\text{Koop}}$ struggles with constraint satisfaction across the full length of all training runs.
@@
Figure 8: Comparison of closed-loop prediction results of a PINN ensemble and a vanilla NN ensemble after different numbers of data sampling steps in the real environment (Fig. 3(a), first step). The bounds of each variable (see Tab. 2) are used for scaling to the [-1, 1] range. The red line depicts the true trajectory of $c$ and $T$. Each blue line corresponds to the closed-loop prediction of one ensemble member. Predictions are chained over the full horizon (168 steps) of the episode.
@@ 4 Conclusion
We validate our method using a demand response case study (Mayfrank et al., 2024b ) based on a benchmark CSTR model (Flores-Tlacuahuac and Grossmann, (2006)). The case study involves nonlinear dynamics and hard constraints on system variables; however, due to its small scale it is far less complex than many real-world systems. We compare the performance of our method to that of Koopman eNMPCs trained solely via iterative SI and to NN policies trained via (physics-informed) model-based RL. We find that our method (see Section 2.3) outperforms all other tested approaches when applied to our case study: It reaches higher rewards and does so with better sample efficiency and lower variance between differently seeded training instances, resulting in improved economic performance and constraint satisfaction compared to the benchmark methods.
@@
Recent works aim to improve upon the MBPO algorithm (e.g., Frauenknecht et al., (2024)) and Dyna-style model-based RL in general (e.g., Frauenknecht et al., (2025)). These methods do not require a specific policy architecture. Therefore, they do not interfere with our approach and could be combined with our method for potentially even better performance. Another avenue of possible future research is combining our method with approaches for learning disturbance estimators for offset-free Koopman MPC (e.g., Son et al., (2021, 2022)): Instead of learning modifications to the state bounds, a task-optimal disturbance estimator could be learned to estimate the disagreement between the Koopman model and the PINN ensemble.